# Recursive Language Models

### 31 Dec 2025 v1

@@ Abstract
We study allowing large language models (LLMs) to process arbitrarily long prompts through the lens of inference-time scaling. We propose Recursive Language Models (RLMs), a general inference strategy that treats long prompts as part of an external environment and allows the LLM to programmatically examine, decompose, and recursively call itself over snippets of the prompt. We find that RLMs successfully handle inputs up to two orders of magnitude beyond model context windows and, even for shorter prompts, dramatically outperform the quality of base LLMs and common long-context scaffolds across four diverse long-context tasks, while having comparable (or cheaper) cost per query.
@@ 1 Introduction
Figure 1: A comparison of GPT-5 and a corresponding RLM on three long-context tasks of increasing complexity: S-NIAH, OOLONG, and OOLONG-Pairs. For each task, we scale the input length from $2^{13}$ to $2^{18}$. GPT-5 performance degrades significantly as a function of both input length and task complexity, while the RLM maintains strong performance. Inputs beyond the red region do not fit in GPT-5’s context window of 272K tokens, but the RLM handles them effectively. Additional experiments across other models, methods, and benchmarks are in §2.
@@
Despite rapid progress in reasoning and tool use, modern language models still have limited context lengths and, even within these limits, appear to inevitably exhibit context rot (Hong et al., 2025), the phenomenon illustrated in the left-hand side of Figure 1 where the quality of even frontier models like GPT-5 degrades quickly as context gets longer. Though we expect context lengths to steadily rise through improvements to training, architecture, and infrastructure, we are interested in whether it possible to dramatically scale the context size of general-purpose LLMs by orders of magnitude. This is increasingly urgent as LLMs begin to be widely adopted for long-horizon tasks, in which they must routinely process tens if not hundreds of millions of tokens.
@@
We study this question through the lens of scaling inference-time compute. We draw broad inspiration from out-of-core algorithms, in which data-processing systems with a small but fast main memory can process far larger datasets by cleverly managing how data is fetched into memory. Inference-time methods for dealing with what are in essence long-context problems are very common, though typically task-specific. One general and increasingly popular inference-time approach in this space is context condensation or compaction (Khattab et al., 2021; Smith, 2025; OpenAI, 2025a; Wu et al., 2025), in which the context is repeatedly summarized once it exceeds a length threshold. Unfortunately, compaction is rarely expressive enough for tasks that require dense access to many parts of the prompt, as it presumes in effect that some details that appear early in the prompt can safely be forgotten to make room for new content.
@@
Figure 2: A Recursive Language Model (RLM) treats prompts as part of the environment. It loads the input prompt as a variable inside a Python REPL environment $\mathcal{E}$ and writes code to peek into, decompose, and invoke itself recursively over programmatic snippets of the variable.
@@
We introduce Recursive Language Models (RLMs), a general-purpose inference paradigm for dramatically scaling the effective input and output lengths of modern LLMs. The key insight is that long prompts should not be fed into the neural network (e.g., Transformer) directly but should instead be treated as part of the environment that the LLM can symbolically interact with.
@@
As Figure 2 illustrates, an RLM exposes the same external interface as an LLM: it accepts a string prompt of arbitrary structure and produces a string response. Given a prompt $P$, the RLM initializes a Read-Eval-Print Loop (REPL) programming environment in which $P$ is set as the value of a variable. It then offers the LLM general context about the REPL environment (e.g., the length of the string $P$), and permits it to write code that peeks into and decomposes $P$, and to iteratively observe any side effects from execution. Crucially, RLMs encourage the LLM, in the code it produces, to programmatically construct sub-tasks on which they can invoke themselves recursively.
By treating the prompt as an object in the external environment, this simple design of RLMs tackles a foundational limitation in the many prior approaches (Anthropic, 2025; Sentient, 2025; Schroeder et al., 2025; Sun et al., 2025), which focus on recursive decomposition of the tasks but cannot allow their input to scale beyond the context window of the underlying LLM.
@@
We evaluate RLMs using a frontier closed model (GPT-5; OpenAI 2025c) and a frontier open model (Qwen3-Coder-480B-A35B; Team 2025) across four diverse tasks with varying levels of complexity for deep research (Chen et al., 2025), information aggregation (Bertsch et al., 2025), code repository understanding (Bai et al., 2025), and a synthetic pairwise reasoning task where even frontier models fail catastrophically. We compare RLMs against direct LLM calls as well as context compaction, retrieval tool-use agents, and code-generation agents. We find that RLMs demonstrate extremely strong performance even at the 10M+ token scale, and dramatically outperform all other approaches at long-context processing, in most cases by double-digit percentage gains while maintaining a comparable or lower cost. In particular, as demonstrated in Figure 1 exhibit far less severe degradation for longer contexts and more sophisticated tasks.
@@ 2 Scaling Long Context Tasks
Recent work (Hsieh et al., 2024; Goldman et al., 2025; Hong et al., 2025) has successfully argued that the effective context window of LLMs can often be much shorter than a model’s physical maximum number of tokens. Going further, we hypothesize that the effective context window of an LLM cannot be understood independently of the specific task. That is, more “complex” problems will exhibit degradation at even shorter lengths than simpler ones. Because of this, we must characterize tasks in terms of how their complexity scales with prompt length.
For example, needle-in-a-haystack (NIAH) problems generally keep ‘needles’ constant as prompt length is scaled. As a result, while previous generations of models struggled with NIAH tasks, frontier models can reliably solve these tasks in RULER (Hsieh et al., 2024) even in the 1M+ token settings. Nonetheless, the same models struggle even at shorter lengths on OOLONG (Bertsch et al., 2025), which is a task where the answer depends explicitly on almost every line in the prompt. [^1] This intuition helps explain the patterns seen in Figure 1 earlier: GPT-5 scales effectively on the S-NIAH task, where the needle size is constant despite longer prompts, but shows faster degradation at increasingly shorter context lengths on the linear complexity OOLONG and the quadratic complexity OOLONG-Pairs.
@@ 2.1 Tasks
Grounded in this intuition, we design our empirical evaluation around tasks where we are able to vary not just the lengths of the prompts, but also consider different scaling patterns for problem complexity. We loosely characterize each task by information density, i.e. how much information an agent is required to process to answer the task, and how this scales with different input sizes.
@@
S-NIAH. Following the single needle-in-the-haystack task in RULER (Hsieh et al., 2024), we consider a set of 50 single needle-in-the-haystack tasks that require finding a specific phrase or number in a large set of unrelated text. These tasks require finding a single answer regardless of input size, and as a result scale roughly constant in processing costs with respect to input length.
@@
BrowseComp-Plus (1K documents) (Chen et al., 2025). A multi-hop question-answering benchmark for DeepResearch (OpenAI, 2025b) questions that requires reasoning over multiple different documents. The benchmark provides a verified offline corpus of 100K documents that is guaranteed to contain gold, evidence, and hard negative documents for each task. Following Sun et al. (2025), we use 150 randomly sampled tasks as our evaluation set; we provide $1000$ randomly chosen documents to the model or agent, in which the gold and evidence documents are guaranteed to exist. We report the percentage of correct answers. The answer to each task requires piecing together information from several documents, making these tasks more complicated than S-NIAH despite also requiring a constant number of documents to answer.
@@
OOLONG (Bertsch et al., 2025). A long reasoning benchmark that requires examining and transforming chunks of the input semantically, then aggregating these chunks to form a final answer. We report scoring based on the original paper, which scores numerical answers as $\texttt{score}(\hat{y})=0.75^{|y-\hat{y}|}$ and other answers as exact match. We focus specifically on the trec_coarse split, which is a set of $50$ tasks over a dataset of questions with semantic labels. Each task requires using nearly all entries of the dataset, and therefore scales linearly in processing costs relative to the input length.
@@
OOLONG-Pairs. We manually modify the trec_coarse split of OOLONG to include $20$ new queries that specifically require aggregating pairs of chunks to construct the final answer. In Appendix E.1, we explicitly provide all queries in this benchmark. We report F1 scores over the answer. Each task requires using nearly all pairs of entries of the dataset, and therefore scales quadratically in processing costs relative to the input length.
LongBench-v2 CodeQA (Bai et al., 2025). A multi-choice code repository understanding split from LongBench-v2 that is challenging for modern frontier models. We report the score as the percentage of correct answers. Each task requires reasoning over a fixed number of files in a codebase to find the right answer.
2.2 Methods and Baselines
We compare RLMs against other commonly used task-agnostic methods. For each of the following methods, we use two contemporary LMs, GPT-5 with medium reasoning (OpenAI, 2025c) and default sampling parameters and Qwen3-Coder-480B-A35B (Yang et al., 2025) using the sampling parameters described in Team (2025), chosen to provide results for a commercial and open frontier model respectively. For Qwen3-Coder, we compute costs based on the Fireworks provider (Fireworks, 2025). In addition to evaluating the base model on all tasks, we also evaluate the following methods and baselines:
RLM with REPL. We implement an RLM that loads its context as a string in the memory of a Python REPL environment. The REPL environment also loads in a module that allows it to query a sub-LM inside the environment. The system prompt is fixed across all experiments (see Appendix D). For the GPT-5 experiments, we use GPT-5-mini for the recursive LMs and GPT-5 for the root LM, as we found this choice to strike a powerful tradeoff between the capabilities of RLMs and the cost of the recursive calls.
RLM with REPL, no sub-calls. We provide an ablation of our method. In it, the REPL environment loads in the context, but is not able to use sub-LM calls. In this setting, the LM can still interact with its context in a REPL environment before providing a final answer.
Summary agent. Following Sun et al. (2025); Wu et al. (2025); Yu et al. (2025), we consider an iterative agent that invokes a summary of the context as it is filled. For example, given a corpus of documents, it will iteratively view the documents and summarize when full. In cases where the provided context exceeds the model window, the agent will chunks the input to fit within the model context window and invoke the same strategy over these chunks. For GPT-5, due to the extremely high cost of handling large token inputs, we use GPT-5-nano for compaction and GPT-5 to provide the final answer.
CodeAct (+ BM25). We compare directly to a CodeAct (Wang et al., 2024) agent that can execute code inside of a ReAct (Yao et al., 2023) loop. Unlike an RLM, it does not offload its prompt to the code environment, and instead provides it directly to the LM. Furthermore, following Jimenez et al. (2024); Chen et al. (2025), we equip this agent with a BM25 (Robertson and Zaragoza, 2009) retriever that indexes the input context for tasks where this is appropriate.
3 Results and Discussion
We focus our main experiments in Table 1 on the benchmarks described in §2.1. Furthermore, we explore how frontier model and RLM performance degrades as input contexts grow in Figure 1.
@@ 3 Results and Discussion
Table 1: Performance comparison of different methods across long-context benchmarks of varying complexity. In gray is the average API cost $\pm$ the standard deviation of each method on each task. indicates runs where the method ran into input context limits.
@@
Qwen3-Coder-480B
Base Model 20.00 (\$0.13 $\pm$ \$0.08) 0.00 (N/A) $\pm$ (N/A) 36.00 (\$0.06 $\pm$ \$0.00) 0.06 (\$0.05 $\pm$ \$0.01)
CodeAct (+ BM25) 24.00 (\$0.17 $\pm$ \$0.08) 12.66 (\$0.39 $\pm$ \$0.50) 38.00 (\$1.51 $\pm$ \$1.09) 0.28 (\$1.54 $\pm$ \$0.35)
Summary agent 50.00 (\$1.26 $\pm$ \$1.50) 38.00 (\$8.98 $\pm$ \$2.12) 44.06 (\$0.15 $\pm$ \$0.01) 0.31 (\$0.05 $\pm$ \$0.00)
RLM 56.00 (\$0.92 $\pm$ \$1.23) 44.66 (\$0.84 $\pm$ \$0.63) 48.00 (\$0.61 $\pm$ \$0.49) 23.11 (\$1.02 $\pm$ \$0.52)
RLM (no sub-calls) 66.00 (\$0.18 $\pm$ \$0.58) 46.00 (\$0.82 $\pm$ \$0.69) 43.50 (\$0.32 $\pm$ \$0.13) 17.34 (\$1.77 $\pm$ \$1.23)
GPT-5
Base Model 24.00 (\$0.13 $\pm$ \$0.07) 0.00 (N/A) $\pm$ (N/A) 44.00 (\$0.14 $\pm$ \$0.02) 0.04 (\$0.16 $\pm$ \$0.10)
CodeAct (+ BM25) 22.00 (\$0.06 $\pm$ \$0.08) 51.00 (\$0.71 $\pm$ \$1.20) 38.00 (\$0.61 $\pm$ \$1.06) 24.67 (\$0.75 $\pm$ \$0.43)
Summary agent 58.00 (\$1.31 $\pm$ \$1.46) 70.47 (\$0.57 $\pm$ \$0.10) 46.00 (\$0.13 $\pm$ \$0.01) 0.01 (\$0.13 $\pm$ \$0.09)
RLM 62.00 (\$0.11 $\pm$ \$0.10) 91.33 (\$0.99 $\pm$ \$1.22) 56.50 (\$0.43 $\pm$ \$0.85) 58.00 (\$0.33 $\pm$ \$0.20)
RLM (no sub-calls) 58.00 (\$0.18 $\pm$ \$0.56) 88.00 (\$0.44 $\pm$ \$0.90) 36.00 (\$0.37 $\pm$ \$0.42) 43.93 (\$0.69 $\pm$ \$1.16)

### 28 Jan 2026 v2

@@ Abstract
We study allowing large language models (LLMs) to process arbitrarily long prompts through the lens of inference-time scaling. We propose Recursive Language Models (RLMs), a general inference paradigm that treats long prompts as part of an external environment and allows the LLM to programmatically examine, decompose, and recursively call itself over snippets of the prompt. We find that RLMs can successfully process inputs up to two orders of magnitude beyond model context windows and, even for shorter prompts, dramatically outperform the quality of vanilla frontier LLMs and common long-context scaffolds across four diverse long-context tasks while having comparable cost. At a small scale, we post-train the first natively recursive language model. Our model, RLM-Qwen3-8B, outperforms the underlying Qwen3-8B model by $28.3\%$ on average and even approaches the quality of vanilla GPT-5 on three long-context tasks. Code is available at https://github.com/alexzhang13/rlm.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: A comparison of GPT-5 and a corresponding RLM using GPT-5 on three long-context tasks of increasing complexity: S-NIAH, OOLONG, and OOLONG-Pairs. For each task, we scale the input length from $2^{13}$ to $2^{18}$. GPT-5 performance degrades significantly as a function of both input length and task complexity, while the RLM maintains strong performance. Inputs beyond the red region do not fit in GPT-5’s context window of 272K tokens, but the RLM handles them effectively. Additional experiments across other models and benchmarks are in §3.
@@
Frontier reasoning models have limited context windows and, even within their limits, tend to exhibit context rot (Hong et al., 2025), a phenomenon illustrated in Figure 1 where quality degrades steeply as prompts get longer. Though we expect context lengths to steadily rise through improvements to training, architecture, and infrastructure, we are interested in whether it is possible to scale the context size of general-purpose LLMs by orders of magnitude. This is increasingly urgent as LLMs begin to be widely adopted for long-horizon tasks, in which they must routinely process tens if not hundreds of millions of tokens.
@@
We study this question through the lens of scaling inference-time compute. We are inspired by the way that reasoning models have become the fundamental interface to LLMs, resulting not only in empirical gains but also additional theoretical expressive power (Merrill and Sabharwal, 2024) compared to vanilla Transformers. Though most inference-time methods for dealing with long context are task-specific (Wu et al., 2021; Chang et al., 2024), the most popular general approach is context condensation or compaction (Khattab et al., 2021; Smith, 2025; OpenAI, 2025a; Wu et al., 2025), where context from user requests or agent trajectories is repeatedly summarized once it exceeds a length threshold. Unfortunately, compaction is rarely expressive enough for tasks that require dense access throughout the prompt. It presumes that some details that appear early in the prompt can safely be forgotten to make room for new content.
@@
Figure 2: A Recursive Language Model (RLM) treats prompts as part of the environment. It loads the input prompt as a variable inside a REPL environment $\mathcal{E}$ and writes code to peek into, decompose, and invoke itself recursively over programmatic snippets of the variable.
@@
We introduce Recursive Language Models (RLMs), a general-purpose inference paradigm for dramatically scaling the effective input and output lengths of LLMs. The key insight is that arbitrarily long user prompts should not be fed into the neural network (e.g., Transformer) directly but should instead be treated as part of the environment that the LLM is tasked to symbolically and recursively interact with.
@@
As Figure 2 shows, an RLM exposes the same external interface as an LLM or a reasoning model: it accepts a string prompt of arbitrary structure and produces a string response. Given a prompt $P$, the RLM initializes a Read-Eval-Print Loop (REPL) programming environment in which $P$ is set as the value of a variable. It then offers the LLM general context about the REPL environment (e.g., the length of the string $P$), and permits it to write code that peeks into and decomposes $P$, and to iteratively observe any side effects from execution. Crucially, RLMs encourage the LLM to understand, transform, and execute the input prompt by writing symbolic programs that invoke the LLM itself on as many slices of the input as necessary.
By treating the prompt itself as an external object and enabling symbolic recursion, RLMs tackle limitations of expressive power in recent work on coding agents, retrieval agents, and sub-agent delegation. In particular, prior coding agents and retrieval agents treat some designated external data source (e.g., a filesystem or a corpus of search documents) as an environment for fetching snippets. However, they can only fill up the underlying LLM’s context window with snippets before breaking down. Similarly, prior self-delegation approaches (Anthropic, 2025; Sentient AI, 2025; Schroeder et al., 2025; Sun et al., 2025) allow LLMs to invoke themselves as sub-agents. However, they are handicapped by the underlying LLM’s limited output lengths because they are designed to verbalize sub-calls autoregressively rather than producing them programmatically.
@@
We evaluate RLMs using a frontier closed model (GPT-5; Singh et al. 2025) and a frontier open model (Qwen3-Coder-480B-A35B; Qwen Team 2025b) across four tasks with varying levels of complexity: deep research (Chen et al., 2025), information aggregation (Bertsch et al., 2025), code repository understanding (Bai et al., 2025), and a synthetic pairwise reasoning task where even frontier models fail catastrophically. We compare RLMs against direct LLM calls as well as context compaction, retrieval tool-use agents, and code-generation agents.
We find that RLMs demonstrate extremely strong performance even at the 10M+ token scale, and substantially outperform all other approaches at long-context processing, in many cases by double-digit percentage gains while maintaining comparable cost. In particular, as demonstrated in Figure 1, RLMs exhibit far less severe degradation for longer contexts and more sophisticated tasks.
Finally, at a small scale, we post-train the first natively recursive language model, demonstrating that RLMs can be improved quickly with little additional training. While a small open model (Qwen3-8B; Yang et al. 2025) struggles to solve long context tasks even in an RLM scaffold, our simple general-purpose training recipe uses only 1,000 samples from unrelated domains to improve its performance by a median of $28.3\%$ across the four evaluation tasks.
2 Recursive Language Models
Given a base neural language model $\mathcal{M}$ with maximum context size $K$, a Recursive Language Model (RLM) is an inference-time scaffold around $\mathcal{M}$ that treats the user prompt as part of the environment without giving up the ability to densely process its content through different calls to $\mathcal{M}$. Given an arbitrary-length prompt string $P\in\Sigma^{\star}$, an RLM interacts with a persistent external environment $\mathcal{E}$ and returns a response string $Y\in\Sigma^{\star}$ (Figure 2). We would like effectively unbounded input tokens ($|P|\gg K$), unbounded output tokens, and an unbounded semantic horizon, e.g. the ability to do $\Omega(|P|)$ or $\Omega(|P|^{2})$ semantic work.
Algorithm 1 describes how an RLM achieves this. Given a prompt $P$, the RLM initializes a persistent REPL programming environment with a variable containing the user prompt as a string and a function for invoking a sub-RLM with a new prompt. Then, it starts the RLM loop. In the first iteration, the algorithm invokes the root neural model $\mathcal{M}$ with only (constant-size) metadata about the user prompt, like its length, a short prefix, and how to access parts of it.
The root is instructed via prompting (Appendix C) and/or fine-tuning (Appendix A) to operate like an RLM: that is, to generate code that helps it understand and transform its parts of its prompt $P$, and to build up intermediate values and the final response into new variables, potentially by invoking the sub-RLM within loops. In Section 4, we find that existing LLMs can be prompted to do this and that training an 8B model to be natively recursive is promising.
Each iteration of the RLM loop executes code in the REPL, updates REPL state (intermediate variables), and collects in stdout any printed text. Only (constant-size) metadata about stdout, like a short prefix and length, is appended to $\mathcal{M}$’s history for the next iteration. [^1] This is key: it forces $\mathcal{M}$ to rely on variables and sub-calls to manage long strings instead of polluting its window. In principle, if we trim each turn to $c$ tokens, we will have at most $K/c$ root iterations, each of which can launch arbitrarily many sub-calls. This is not a fundamental limitation, e.g. one could move the root horizon itself into a variable, but we typically want to limit the iterations at any level of recursion irrespective. Once the RLM sets the variable Final inside the REPL, iteration stops and the value in Final is returned as the response.
RLMs make three simple design choices that are missing from existing scaffolds. To highlight these, we include Algorithm 2 to illustrate a deceptively “similar” algorithm that is far less expressive. Both algorithms support some notion of sub-calls, external objects, and code execution, but they differ in terms of where the prompt and intermediate values live and where recursion occurs.
Algorithm 1 A recursive language model, around LLM $\mathcal{M}$
Input: prompt $P$
Output: response $Y$
state $\leftarrow$ InitREPL(prompt=P)
state $\leftarrow$ AddFunction(state,  sub_RLMM)
hist $\leftarrow[\texttt{Metadata(state)}]$
while True do
 code $\leftarrow$ LLMM(hist)  (state, stdout) $\leftarrow$ REPL(state, code)hist $\leftarrow$ hist $\|\,$ code $\|\,$ Metadata(stdout)if state[Final] is set then
    return state[Final]
Algorithm 2 Alternate scaffold with standard (poor) design choices for prompts, sub-calls, and code execution
Input: prompt $P$
Output: response $Y$
actions $\leftarrow\{\texttt{Finish},\,\texttt{Exec},\,\texttt{Search},\,{\color[rgb]{0.78515625,0,0.78515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.78515625,0,0.78515625}\texttt{sub\_LLM}}_{\mathcal{M}}\}$
hist $\leftarrow[\texttt{Metadata(actions)},\,P]$
// Flaw #1while True do
   (action, val) $\leftarrow$ LLMM(hist)if action is Finish then
    return val
    // Flaw #2
 out $\leftarrow$ RUN(action, val)
 // Flaw #3hist $\leftarrow$ hist $\|$ (action, val, out)  if Tok(hist) > K then
    hist $\leftarrow$ Compact(hist)
First, an RLM must give the underlying LLM $\mathcal{M}$ a symbolic handle to the user prompt $P$, so the model can manipulate it without copying text into the root context window. Instead, ineffective Algorithm 2 starts by putting the user prompt $P$ into the LLM context window (hist) and thus inherits the window limitations of $\mathcal{M}$ and falls back to heuristics like context compaction. Even though the scaffold can access external data with, say, a Search action or filesystem access, it is fatally bounded with respect to user input.
Second, ineffective Algorithm 2 asks $\mathcal{M}$ to autoregressively generate the output directly, via a Finish action. This may seem innocuous, but it means that it also cannot generate longer outputs than the context window of $\mathcal{M}$ permits.
Third, and perhaps most importantly, an RLM requires symbolic recursion. That is, code running inside $\mathcal{E}$ must be able to invoke $\mathcal{M}$ on programmatically constructed transformations of $P$ (e.g., inside arbitrarily large loops), storing intermediate results symbolically. Though Algorithm 2 includes both a code execution action and a “sub-LLM” action separately, it is not able to invoke the sub-LLM programmatically and hence can only delegate a few explicitly verbalized tasks rather than writing short programs that can, say, loop over slices of the prompt and launch $\Omega(|P|)$ or even $\Omega(|P|^{2})$ processes to understand or transform all parts of $P$.
@@ 3 Scaling Long Context Tasks
We hypothesize that the effective context window (Hsieh et al., 2024; Goldman et al., 2025; Hong et al., 2025) of an LLM cannot be understood independently of the specific task. That is, more “complex” problems will exhibit degradation at even shorter lengths than simpler ones. Because of this, we must characterize tasks in terms of how their complexity scales with prompt length.
For example, needle-in-a-haystack (NIAH) problems generally keep ‘needles’ constant as prompt length is scaled. As a result, frontier models can now reliably solve these tasks in RULER (Hsieh et al., 2024) in the 1M+ token settings but struggle at far shorter lengths on OOLONG (Bertsch et al., 2025), a task where the answer depends explicitly on almost every line in the prompt. [^2] This helps explain the patterns seen in Figure 1 earlier: GPT-5 scales effectively on the S-NIAH task, where the needle size is constant despite longer prompts, but shows faster degradation at increasingly shorter context lengths on the linear-complexity OOLONG and the quadratic-complexity OOLONG-Pairs.
@@ 3.1 Tasks
We design our evaluation around tasks where we can vary the lengths of the prompts, so we can consider problems whose difficulties scale differently with context length.
@@
S-NIAH. Following the single needle-in-the-haystack task in RULER (Hsieh et al., 2024), we consider a set of 50 single tasks that require finding a specific phrase or number in a large set of unrelated text. Here, the information being sought scales as $O(1)$ with respect to input length.
@@
BrowseComp-Plus (1K documents) (Chen et al., 2025). A multi-hop question-answering benchmark for DeepResearch (OpenAI, 2025b) questions that requires reasoning over multiple different documents. The benchmark provides a verified offline corpus that is guaranteed to contain gold, evidence, and hard negative documents for each question. Following Sun et al. (2025), we use 150 randomly sampled instances as our evaluation set; we provide $1000$ randomly chosen documents as input, in which the gold and evidence documents are guaranteed to exist. We report the percentage of correct answers. The answer to each task requires piecing together information from several documents, making this harder than S-NIAH despite also requiring a constant number of documents.
@@
OOLONG (Bertsch et al., 2025). A long reasoning benchmark that requires transforming chunks of the input semantically, then aggregating these chunks to form a final answer. We report scoring based on the original paper, which scores numerical answers as $\texttt{score}(\hat{y})=0.75^{|y-\hat{y}|}$ and other answers as exact match. We focus specifically on the trec_coarse split, a set of $50$ tasks over a dataset of questions with semantic labels. Each task requires using nearly all entries of the dataset, and therefore scales linearly in processing complexity relative to the input length.
@@
OOLONG-Pairs. We modify the trec_coarse split of OOLONG to include $20$ new queries that specifically require aggregating pairs of chunks to construct the final answer. We report F1 scores over the answer. Each task requires using nearly all pairs of entries of the dataset, and therefore requires processing quadratically-many items relative to the input length. In Appendix D.1, we provide all queries in this benchmark.
@@
Table 1: Performance comparison of different methods across long-context benchmarks of varying complexity. In gray is the average API cost $\pm$ the standard deviation of each method on each task. indicates runs where a method (sometimes) ran into input context limits. Provider costs were computed under OpenAI for GPT-5 and Fireworks for other models. Non-zero scores are rounded to at least $0.1$.
@@
GPT-5 (with RLM sub-calls to GPT-5-mini)
Base Model 24.0 (\$0.13 $\pm$ \$0.07) 0.0 (N/A) $\pm$ (N/A) 44.0 (\$0.14 $\pm$ \$0.02) 0.1 (\$0.16 $\pm$ \$0.10)
CodeAct (+ BM25) 22.0 (\$0.06 $\pm$ \$0.08) 51.0 (\$0.71 $\pm$ \$1.20) 38.0 (\$0.61 $\pm$ \$1.06) 24.7 (\$0.75 $\pm$ \$0.43)
CodeAct (+ sub-calls) 24.0 (\$0.06 $\pm$ \$0.08) 0.0 (N/A) $\pm$ (N/A) 40.0 (\$0.85 $\pm$ \$1.27) 28.4 (\$1.11 $\pm$ \$0.62)
Summary agent 58.0 (\$1.31 $\pm$ \$1.46) 70.5 (\$0.57 $\pm$ \$0.10) 46.0 (\$0.13 $\pm$ \$0.01) 0.1 (\$0.13 $\pm$ \$0.09)
RLM 62.0 (\$0.11 $\pm$ \$0.10) 91.3 (\$0.99 $\pm$ \$1.22) 56.5 (\$0.43 $\pm$ \$0.85) 58.0 (\$0.33 $\pm$ \$0.20)
RLM (no sub-calls) 58.0 (\$0.18 $\pm$ \$0.56) 88.0 (\$0.44 $\pm$ \$0.90) 36.0 (\$0.37 $\pm$ \$0.42) 43.9 (\$0.69 $\pm$ \$1.16)
Qwen3-Coder-480B-A35B
Base Model 20.0 (\$0.13 $\pm$ \$0.08) 0.0 (N/A) $\pm$ (N/A) 36.0 (\$0.06 $\pm$ \$0.00) 0.1 (\$0.05 $\pm$ \$0.01)
CodeAct (+ BM25) 24.0 (\$0.17 $\pm$ \$0.08) 12.7 (\$0.39 $\pm$ \$0.50) 38.0 (\$1.51 $\pm$ \$1.09) 0.3 (\$1.54 $\pm$ \$0.35)
CodeAct (+ sub-calls) 26.0 (\$0.28 $\pm$ \$0.30) 0.0 (N/A) $\pm$ (N/A) 32.0 (\$1.83 $\pm$ \$1.14) 0.1 (\$1.49 $\pm$ \$0.46)
Summary agent 50.0 (\$1.26 $\pm$ \$1.50) 38.0 (\$8.98 $\pm$ \$2.12) 44.1 (\$0.15 $\pm$ \$0.01) 0.31 (\$0.05 $\pm$ \$0.00)
RLM 56.0 (\$0.92 $\pm$ \$1.23) 44.7 (\$0.84 $\pm$ \$0.63) 48.0 (\$0.61 $\pm$ \$0.49) 23.1 (\$1.02 $\pm$ \$0.52)
RLM (no sub-calls) 66.0 (\$0.18 $\pm$ \$0.58) 46.0 (\$0.82 $\pm$ \$0.69) 43.5 (\$0.32 $\pm$ \$0.13) 17.3 (\$1.77 $\pm$ \$1.23)
Qwen3-8B
Base Model 4.0 (\$0.01 $\pm$ \$0.00) 0.0 (N/A) $\pm$ (N/A) 0.0 (N/A) $\pm$ (N/A) 0.1 (\$0.01 $\pm$ \$0.00)
RLM 26.0 (\$0.04 $\pm$ \$0.13) 2.0 (\$0.03 $\pm$ \$0.06) 24.0 (\$0.19 $\pm$ \$0.26) 4.3 (\$0.05 $\pm$ \$0.05)
RLM (fine-tuned) 32.0 (\$0.02 $\pm$ \$0.02) 14.0 (\$0.01 $\pm$ \$0.03) 32.0 (\$0.04 $\pm$ \$0.09) 5.2 (\$0.02 $\pm$ \$0.02)
LongBench-v2 CodeQA (Bai et al., 2025). A multi-choice code repository understanding split from LongBench-v2 that is challenging for modern frontier models. We report the score as the percentage of correct answers. Each instance requires reasoning over a fixed number of files in a codebase to find the right answer.
3.2 Methods and Baselines
We compare RLMs against commonly used task-agnostic inference methods, using two modern LMs, GPT-5 with medium reasoning (Singh et al., 2025) and default sampling parameters, and Qwen3-Coder-480B-A35B (Yang et al., 2025) using the sampling parameters described in Qwen Team (2025b). For Qwen3-Coder-480B-A35B, we compute costs based on the compute provider Fireworks (Fireworks AI, 2025). In addition to evaluating the base model on all tasks, we also evaluate the following methods and baselines:
CodeAct (+ BM25). We compare directly to a CodeAct (Wang et al., 2024) agent that can execute code inside of a ReAct (Yao et al., 2023) loop. Unlike an RLM, CodeAct does not offload the user prompt to the code environment, and instead provides it directly to the LM. Furthermore, following Jimenez et al. (2024); Chen et al. (2025), we equip this agent with a BM25 (Robertson and Zaragoza, 2009) retriever that indexes the input context for tasks where a retriever is appropriate.
CodeAct with sub-calls. To specifically ablate offloading the context as a variable in the REPL, we evaluate a CodeAct (Wang et al., 2024) baseline with the ability to invoke sub-LM calls. Compared to RLMs, this method loads the context directly into the model.
Summary agent. Following Sun et al. (2025); Wu et al. (2025); Yu et al. (2025), we consider an iterative agent that compacts the context as it is filled. For example, given a corpus of documents, it will iteratively accumulate the documents and summarize when full. In cases where a single document exceeds the model window, the agent will chunk it to fit within the model context window and invoke the same strategy over these chunks. For the GPT-5 experiments, due to the extremely high cost of applying this strategy to millions of tokens, we use GPT-5-nano for compaction and GPT-5 to provide the final answer.
RLM with REPL. We implement an RLM with a Python REPL environment, which loads a module for querying a sub-LM and uses a system prompt presented in Appendix C. For the GPT-5 experiments, we use GPT-5-mini for the recursive LMs and GPT-5 for the root LM, as we found this choice to strike a good balance between the capabilities of RLMs and the cost of the recursive calls. We notate a RLM using a model as RLM(model), e.g. RLM(GPT-5).
RLM with REPL, no sub-calls. We provide an ablation of our method, in which the prompt is loaded in a REPL environment without the ability to invoke sub-LM calls.
Finetuning. To create RLM-Qwen3-8B, we finetune Qwen3-8B on 1,000 filtered trajectories of Qwen3-Coder-480B-A35B as an RLM with Qwen3-8B sub-calls on LongBenchPro (Chen et al., 2026) tasks. We use sampling parameters described in Qwen Team (2025a), and evaluate the fine-tuned RLM-Qwen3-8B as an RLM on our long context tasks. The key insight for training is that being an effective sub-call model is roughly similar to being a general purpose reasoning model, so we can make the training much more tractable (and seemingly short-horizon) at small scale by focusing on improving the root model’s ability to manipulate the REPL and to launch recursive calls. We provide more training details in Appendix A.
4 Results and Discussion
Table 1 reports our main results. We additionally explore how vanilla frontier model performance and RLM performance degrades as input contexts grow in Figure 1.